A solid body starts rotating about a stationary axis with an angular acceleration
a =(0.02)t rad/second square, here t is in seconds. How soon after the beginning of rotation, will the total acceleration vector of an arbitrary point of the body form an angle ?=60, with its velocity vector?

Grade:11

1 Answers

Prudhvi teja
13 years ago

Dear pavan

To get angle 60 with velocity vector

ar/an= tan60

=>w^2r/ran = 3^0.5

=> w^2 = an*3^0.5

=>w= 0.186

t=w/an(initial w=0)

=9.3sec

Please feel free to post as many doubts on our discussion forum as you can.we will get you the answer and detailed  solution very  quickly.

All the best.

Regards,
Askiitians Experts
Prudhvi Teja

 

 


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitia